Foreign-Earned Income Exclusion. Fred, a U.S. citizen, arrives in Country K on July 15,
2016, and proceeds to a construction site in its oil fields. Once there, he moves into employer-
provided housing where he is required to reside. Except for brief periods of local
travel and the months of July and August 2017, when he is on vacation in the United
States, he remains at the site until his departure on December 1, 2018. He provides no services
while in the United States. Fred earns $10,000 per month in salary and allowances
while employed overseas. In addition, while in Country K, he receives meals and lodging
valued at $1,750 per month. What is the amount of Fred’s Sec. 911 exclusions for 2016
through 2018?
Answer #1
Foreign-Earned Income Exclusion
A U.S. citizen or a resident alien of the United States is eligible for the deduction on foreign earned income.
The maximum deduction for the Foreign earned incomes are:
Year Maximum Deduction
2015 92900
2016 95100
2017 97600
2018
99200
2019 100800
Amount of deduction for the year 2016   
Actual salary received = 10000*12 = 120000
Maximum deduction restricted to $95100
Amount of deduction for the year 2018   
Actual salary received = 10000*12 = 120000
Maximum deduction restricted to $99200

In addition to the above exclusion, exclusion for meals and lodging also available subject to the following conditions

1. The meals are furnished:

a. On the business premises of your employer, and

b. For the convenience of your employer.

2. The lodging is furnished:

a. On the business premises of your employer,

b. For the convenience of your employer, and

c. As a condition of your employment.

It is assumed that the assessee has satisfied all the conditions and he is eligible for the entire meals and lodging exclusion also.